Fifteen identical particles have various speeds. One has a speed of 3.00 m/s, two have a speed of 4.00 m/s, three have a speed of 6.00 m/s, four have a speed of 6.00 m/s, three have a speed of 8.00 m/s and two have a speed of 10.0 m/s. Find (a) the average speed, (b) the rms speed, and (c) the most probable speed of these particles.

Question 17 options:

a

(a) 6.17 m/s; (b) 8.76 m/s; (c) 6.00 m/s

b

(a) 6.47 m/s; (b) 6.76 m/s; (c) 7.00 m/s

c

(a) 6.47 m/s; (b) 6.76 m/s; (c) 10.0 m/s

d

(a) 6.47 m/s; (b) 6.76 m/s; (c) 6.00 m/s

e

(a) 6.80 m/s; (b) 7.41 m/s; (c) 7.00 m/s
Answer #1

n ) -> Given 15 identical partides have various speeds. 3,4,4,6,6,6,6,6,6,6,8,8,8, 10, 10 m/s a) The average speed N=15 (Tote2 GI مع 1X (3)+2x(4)+7X16)+ 3x (8)2+ 2X (10)? = IS 9+32+252 +192 +200 685 15 15 ūd = 45.66m/s =) Usims - J45.66 =) Urms = 6.7